# ModHiFi: Identifying High Fidelity predictive components for Model Modification

### v1

@@ 2.3 Related Work
Vision Model Modification While structured pruning is well-established for CNNs and ViTs [hoefler2021sparsity, fang2024isomorphic, yu2022width, fang2024isomorphic, yu2022unified], and classwise unlearning has seen recent progress [jia2023model, fan2023salun], these tasks are typically treated in isolation. Crucially, prior methods for jointly addressing these problems rely heavily on access to labeled data [murti2024discedit]. Our work provides the first unified framework for both pruning and unlearning that operates effectively using only unlabeled synthetic data.
LLM Modification Efficiency in LLMs is primarily addressed via structured pruning [ma2023llm, men2024shortgpt] or sparsification [ashkboos2024slicegpt]. However, these methods are often architecture-specific and do not extend to unlearning. By validating our method on both LLMs and vision models, we demonstrate a generalized approach to model modification that bridges the gap between these distinct domains.
@@ 3 Which Components Are Important for Modifying Well-Trained Models?
We now address the problem of identifying model components critical to predictive performance. We introduce Subset Fidelity, a metric quantifying the local reconstructive capacity of component groups, and High-Fidelity (HiFi) components. We show theoretically that maximizing local fidelity minimizes a linear upper bound on the global predictive error.
@@ 3.1 High-Fidelity Components and the Subset Fidelity Score
Our objective is to estimate how removing a subset of input contributions affects the model’s output, after optimally compensating for this removal. Directly quantifying this effect is difficult, so we introduce the Subset Fidelity, a measure of how well a subset of components can locally approximate the layer’s output.
@@
$\mathrm{FS}^{l}_{c}(C)\coloneqq\max_{\bm{\delta}^{l}_{c}\in\mathbb{R}^{c_{in}^{l}}}\left(1-\frac{\mathbb{E}\left[\|\bm{Y}^{l}_{c}(\mathrm{X})-\sum_{i\in C}\delta^{l}_{ci}\bm{A}^{l}_{ci}(\mathrm{X})\right\|^{2}]}{\mathbb{E}\left[\|\bm{Y}^{l}_{c}(\mathrm{X})\|^{2}\right]}\right),$ (1)
@@
[Properties of Subset Fidelity]lemmaFSProp For any subset $C\subseteq[c^{l}_{in}]$ in layer $l$,
  1. (1)
    Boundedness: $0\leq\mathrm{FS}^{l}_{c}(C)\leq 1$.
  2. (2)
    Monotonicity: If $D\subseteq C$, then $\mathrm{FS}^{l}_{c}(D)\leq\mathrm{FS}^{l}_{c}(C)$.
A larger Subset Fidelity indicates that the subset more effectively reconstructs the output, thereby, reducing the error of approximating the sum of components with components from a subset. Section˜3.1 implies two key insights: (1) fidelity serves as a principled measure of component importance, and (2) monotonicity suggests that greedy selection strategies may be effective.
@@
Equation˜1 is a generalizes the formulation of elhalabi2022dataefficient. In this work, we focus only on the case where the subset fidelities are measured with the expected squared difference. We leave to future work an exploration of other possible measures of distributional similarity.
To capture the tradeoff between the size of a subset and it’s fidelity, we define HiFi Sets.
@@
A simple algorithm for identifying a $(k,\eta)$-HiFi set is to solve Equation˜k-MFS for the given $k$ and check whether its fidelity exceeds $\eta$. If it does not, no such $(k,\eta)$-HiFi set exists.
Existence of small HiFi components
Before proceeding to our theoretical analysis, we empirically verify whether small HiFi sets actually exist in standard models. Our experiments in Section˜5.2 empirically establish the existence of a small subset of components that can achieve high fidelity. Figure˜1 indicates a sample of the results indicating that fewer than 20% of components can achieve high fidelity ($\geq 0.8$). Moreover, in Section˜5.3, we validate the effectiveness of HiFi components with the model’s predictive performance.
@@
Figure 1: Monte Carlo estimation of Equation˜k-MFS across selected layers of various models. The x-axis indicates subset size $k$, and the y-axis the maximum fidelity found across random samples.
@@ 3.2 Local Distributional Measures of Component Importance
We now show that the influence of a component on its immediate layer output provides a tractable proxy for its overall effect on model predictions.
@@
Consider a network $\mathrm{N}_{\theta}$ as defined in Section˜2.1. Let $\bm{M}^{l}$ be a mask modifying parameters at layer $l$, and let $\bm{m}^{l}_{c}$ be the mask vector for output channel $c$. Assume there exist scalars $r^{\ell}>0$ for all layers $\ell>l$ such that $\|\bm{\Phi}^{\ell}_{c}(\mathrm{X})\|_{F}\geq r^{\ell}$ almost surely. Then, there exist architecture-dependent constants $\{C_{l}\}$ such that
$\mathbb{E}\!\left[\|\mathrm{N}_{\theta}(\mathrm{X})-\mathrm{N}_{\theta\odot\bm{M}^{l}}(\mathrm{X})\|^{2}\right]\leq C_{l}^{2}\sum_{c=1}^{c_{out}^{l}}\left((\bm{1}_{c_{in}^{l}}-\bm{m}^{l}_{c})^{\top}\bm{Q}^{l}_{c}(\bm{1}_{c_{in}^{l}}-\bm{m}^{l}_{c})\right)$ (2)
where $\bm{Q}^{l}_{c}\in\mathbb{R}^{c_{in}^{l}\times c_{in}^{l}}$ is the component similarity matrix (CSM) for channel $c$, with entries $(\bm{Q}^{l}_{c})_{ij}=\mathbb{E}[\langle\bm{A}^{l}_{ci}(\mathrm{X}),\bm{A}^{l}_{cj}(\mathrm{X})\rangle]$.
@@
See Section˜B.1. The proof relies on the propagation of error through Lipschitz-continuous layers. ∎
Section˜3.2 upper-bounds the global reconstruction error by a linear function of the local reconstruction errors for each channel in layer $l$. This implies that global error grows at most linearly with local error, making local fidelity a practical, architecture-agnostic proxy for component influence. The theorem necessitates that networks discussed in this work under suitable conditions are Lipschitz continuous. While CNNs are known to be Lipschitz continuous [yu2018nisp], transformers are not [qi2023lipsformer]. In Corollary˜B.4, we show that this is not the case for well-trained transformers.
@@
The constant $C_{l}$ quantifies the amplification of local errors through subsequent layers and activations, and is independent of the data distribution. Empirical estimates of $C_{l}$ reported in Section˜C.2 demonstrate the practicality of these constants.
Subset Fidelity for $l_{2}$ reconstruction error
@@
[Compensation and Singleton Fidelity]propositioncorr For the $l_{2}$ reconstruction error, the optimal compensation term $\delta^{\star}_{c}$, which is the value at which the fidelity score is computed according to Equation˜1 for a subset $C$ is given by,
@@
where $\bm{Q}^{l}_{c}$ is the component similarity matrix as in Equation˜2. The singleton fidelity scores are:
@@
Note that solving Equation˜k-MFS exactly is still equivalent to a constrained binary quadratic optimization problem, known to be NP-hard [arora_barak]. Viewing $\bm{Q}^{c}$ as the adjacency matrix of a weighted graph, maximizing Equation˜k-MFS corresponds to identifying a clique of size $k$, the decision version of the Maximum Clique problem. Intuitively, such cliques correspond to groups of components whose joint removal maximally increases the reconstruction error. However, since fidelity is monotonic, a natural heuristic selects the $k$ components with the highest singleton fidelities $s^{l}_{ci}$; we call this strategy: NAIVE.
Optimal Identification of HiFi Sets
We identify conditions under which the NAIVE selection strategy is optimal. A proof is provided in Section˜B.4. {restatable}theoremoptlabel Consider output channel $c$ in the $l^{th}$ layer of a network described in Section˜2.1. Let the $s^{l}_{ci}$ be defined according to Equation˜3 and $S^{l\star}_{c}$ be defined according to Definition˜3.2. Let
$\hat{S}^{l}_{c}=\{i\mid s^{l}_{ci}\geq s_{(k)}\}$
where $s_{(k)}$ is the $k^{th}$ largest value of $\bm{s}^{l}_{c}$. Assuming that that there are no ties, $|\hat{S}^{l}_{c}|=k$. If $\mathbb{E}[\langle\bm{A}^{l}_{ci}(\mathrm{X}),\bm{A}^{l}_{cj}(\mathrm{X})\rangle]=0\ \ \forall i\neq j$, then $\hat{S}^{l}_{c}=S^{l\star}_{c}$.
@@
Although the assumption of uncorrelated features rarely holds exactly in practice, it offers a useful theoretical justification for NAIVE HiFi selection. We demonstrate the practical effectiveness of NAIVE HiFi selection through our experiments in Section˜5.
@@ 4 Modifying Model Behavior using HiFi Sets
We now propose ModHiFi, a unified algorithmic framework for model modification using only distributional access. We apply this framework to two distinct tasks: structured pruning (ModHiFi-P) and classwise unlearning (ModHiFi-U). The central idea is to identify high-fidelity (HiFi) components and then modify them in a targeted manner using a unified algorithmic procedure, Algorithm˜1. The two tasks operate as duals: pruning retains the high-fidelity components necessary for general performance, while unlearning removes the high-fidelity components most discriminative for a specific target class. Additional details including complexity and implementation specifics are provided in Appendix˜D.
@@
1:Estimate Fidelity: Compute singleton scores $\mathbf{s}^{l}$ on $\mathcal{D}$ via Equation˜3.
@@
Structured Pruning
In structured pruning, the objective is to remove entire input channels (or features) that contribute minimally to the model’s predictive performance. In convolutional architectures, we identify and remove input channels across all layers that do not appear in the HiFi sets of any output channel of the residual-coupled layers. For CNNs, pruning is applied to the input channels of convolutional layers. For LLMs, we target the input features of the MLP down-projection matrices ($W^{D}$). After pruning, we compute the optimal compensation term $\delta^{\star}$ (derived in Section˜3.2) using the remaining weights. This step restores the fidelity of the layer output without requiring gradient-based fine-tuning.
@@
For unlearning, the goal is to erase the influence of a specific forget class. To perform unlearning, we first compute HiFi sets using only samples from the class we wish to forget. The components in these sets are then zeroed out, effectively erasing the influence of that class. This causes the model’s predictive performance on the forgotten class to degrade without significantly impacting other classes.
@@
For vision models, the singleton fidelity score $\mathrm{FS}_{c}^{l}(\cdot)$ can be estimated efficiently using distributional access to the input data, i.e. synthetic samples. In practice, for vision models, we estimate the scalar coefficients $\alpha_{ci}^{l}$ directly via batched forward passes on synthetic samples. A large $\alpha_{ci}^{l}$ indicates a high-fidelity component. For LLMs, we develop a tractable Cholesky-based heuristic to estimate the score and provide details in Section˜D.2.
@@ 5 Experiments
    Effectiveness of using HiFi components for machine unlearning using ModHiFi-U. Is it possible to perform machine unlearning, as posed by Jia et al. [jia2023model], without fine-tuning? If yes, how does ModHiFi-U fare against their method?
@@ 5.1 Details of the experimental setup
We conduct experiments on ResNet-50/101 [he2016deep], VGG19 [simonyan2015very], Swin-Transformer [liu2021swintransformerhierarchicalvision] and Llama-2-7B [touvron2023llama], benchmarking against relevant experiments from related literature [ashkboos2024slicegpt, ma2023llm]. For vision tasks, we measure the classification accuracy, and for NLP tasks, we use EleutherAI’s lm-eval-harness [eval-harness].
@@
For CIFAR10/100 [krizhevsky2009learning], we use synthetically generated images as detailed in Section˜C.3. We use Alpaca [alpaca] (a synthetic dataset) and WikiText-2 [merity2016pointersentinelmixturemodels] as calibration data for NLP tasks following related literature [ashkboos2024slicegpt, ma2023llm]. We provide ablations to measure the impact of synthetic data quality in Section˜C.3.3.
@@
We discuss the compute platform, implementation details and hyperparameters used for our experiments in Section˜C.6.
@@ 5.2 Existence of HiFi components: Exploring Q1
This empirical observation indicates that in trained models, there are only a small set of components in each layer that are responsible for the model prediction. This observation aligns with the success of structured pruning algorithms in constructing small subnetworks with high statistical performance.

### v2

@@ 2.3 Related Work
Vision Model Modification While structured pruning is well-established for CNNs and ViTs [28, 12, 93, 12, 95], and classwise unlearning has seen recent progress [33, 11], these tasks are typically treated in isolation. Crucially, prior methods for jointly addressing these problems rely heavily on access to labeled data [55]. Our work presents the first unified framework for both pruning and unlearning, which operates effectively using only unlabeled synthetic data.
LLM Modification Efficiency in LLMs is primarily addressed via structured pruning [48, 49] or sparsification [2]. However, these methods are often architecture-specific and do not extend to unlearning. By validating our method on both LLMs and vision models, we demonstrate a generalized approach to model modification that bridges the gap between these distinct domains.
@@ 3 Which Components Are Important for Modifying Well-Trained Models?
We now address the problem of identifying model components critical to predictive performance. We introduce Subset Fidelity, a metric that quantifies the local reconstructive capacity of component groups and High-Fidelity (HiFi) components. We show theoretically that maximizing local fidelity minimizes a linear upper bound on the global predictive error.
@@ 3.1 High-Fidelity Components and the Subset Fidelity Score
Our objective is to estimate the impact of removing a subset of input contributions on the model’s output, after optimally compensating for this removal. Directly quantifying this effect is difficult, so we introduce the Subset Fidelity, a measure of how well a subset of components can locally approximate the layer’s output.
@@
$\mathrm{FS}^{l}_{c}(C)\coloneqq\max_{\bm{\delta}^{l}_{c}\in\mathbb{R}^{c_{in}^{l}}}\left(1-\frac{\mathbb{E}\left[\|\bm{Y}^{l}_{c}(\mathrm{X})-\sum_{i\in C}\delta^{l}_{ci}\bm{A}^{l}_{ci}(\mathrm{X})\|^{2}\right]}{\mathbb{E}\left[\|\bm{Y}^{l}_{c}(\mathrm{X})\|^{2}\right]}\right),$ (1)
@@
[Properties of Subset Fidelity]lemmaFSProp For any subset $C\subseteq[c^{l}_{in}]$ in layer $l$, (Boundedness) $0\leq\mathrm{FS}^{l}_{c}(C)\leq 1$ and (Monotonicity) If $D\subseteq C$, then $\mathrm{FS}^{l}_{c}(D)\leq\mathrm{FS}^{l}_{c}(C)$.
A larger Subset Fidelity indicates that the subset more effectively reconstructs the output, thereby reducing the error of approximating the sum of components with components from a subset. Section˜3.1 implies two key insights: (1) Fidelity serves as a principled measure of component importance, and (2) Monotonicity suggests that greedy selection strategies may be effective.
@@
Eq.˜1 is a generalizes the formulation of Halabi et al. [25]. In this work, we focus only on the case where the subset fidelities are measured with the expected squared difference. We leave to future work an exploration of other possible measures of distributional similarity.
To capture the tradeoff between the size of a subset and its fidelity, we define HiFi Sets.
@@
A simple algorithm for identifying a $(k,\eta)$-HiFi set is to solve Eq.˜k-MFS for the given $k$ and check whether its fidelity exceeds $\eta$. If it does not, no such $(k,\eta)$-HiFi set exists. Before proceeding to our theoretical analysis, we empirically verify whether small HiFi sets actually exist in standard models. Our experiments in Section˜5.2 empirically establish the existence of a small subset of components that can achieve high fidelity. Moreover, in Section˜5.3, we validate the effectiveness of HiFi components with the model’s predictive performance. Fig.˜1 indicates a sample of the results indicating that fewer than 20% of components can achieve high fidelity ($\geq 0.8$).
@@
Figure 1: Monte Carlo estimation of Eq.˜k-MFS across selected layers of various models. The x-axis indicates subset size $k$, and the y-axis the maximum fidelity found across random samples.
@@ 3.2 Local Distributional Measures of Component Importance
We now show that the influence of a component on its immediate layer output provides a tractable proxy for its overall effect on model predictions. The global error is the expectation of the squared difference in the predictions of a network under a modification.
@@
Consider a network $\mathrm{N}_{\theta}$ as defined in Section˜2.1. Let $\bm{M}^{l}$ be a mask modifying parameters at layer $l$, and let $\bm{m}^{l}_{c}$ be the mask vector for output channel $c$. Assume there exist scalars $r^{\ell}>0$ for all layers $\ell>l$ such that $\|\bm{\Phi}^{\ell}_{c}(\mathrm{X})\|_{F}\geq r^{\ell}$ almost surely. Then,
$\mathbb{E}\!\left[\|\mathrm{N}_{\theta}(\mathrm{X})-\mathrm{N}_{\theta\odot\bm{M}^{l}}(\mathrm{X})\|^{2}\right]\leq\mathcal{O}\left(\sum_{c=1}^{c_{out}^{l}}\mathbb{E}\left[\|\bm{Y}^{l}_{c}(\mathrm{X})-\sum_{i\in C}m^{l}_{ci}\bm{A}^{l}_{ci}(\mathrm{X})\|^{2}\right]\right)$ (2)
@@
The proof relies on the propagation of error through Lipschitz-continuous layers. See Section˜B.1. ∎
Section˜3.2 upper-bounds the global error, given by the left-hand side, by a linear function of the local reconstruction errors for each channel in layer $l$. This implies that global error grows at most linearly with local error, making local fidelity a practical, architecture-agnostic proxy for component influence. The theorem requires that the networks discussed in this work are Lipschitz continuous under suitable conditions. While CNNs are known to be Lipschitz continuous [94], transformers are not [63]. In Corollary˜B.4, we show that this is not the case for well-trained transformers.
@@
The leading constant in the order notation quantifies the amplification of local errors through subsequent layers and activations, and is independent of the data distribution, depending only on the model’s architecture. Empirical estimates of the constant reported in Section˜C.2 demonstrate the practicality of these constants.
Subset Fidelity for Individual Components
@@
[Compensation and Singleton Fidelity]propositioncorr For the $l_{2}$ reconstruction error, the optimal compensation term $\delta^{\star}_{c}$, which is the value at which the fidelity score is computed according to Eq.˜1 for a subset $C$, is given by,
@@
where $\bm{Q}^{l}_{c}\in\mathbb{R}^{c_{in}^{l}\times c_{in}^{l}}$ is the component similarity matrix (CSM) for channel $c$, with entries $(\bm{Q}^{l}_{c})_{ij}=\mathbb{E}[\langle\bm{A}^{l}_{ci}(\mathrm{X}),\bm{A}^{l}_{cj}(\mathrm{X})\rangle]$. The singleton fidelity scores are:
@@
Note that solving Eq.˜k-MFS exactly is still equivalent to a constrained binary quadratic optimization problem, known to be NP-hard [1]. Viewing $\bm{Q}^{c}$ as the adjacency matrix of a weighted graph, maximizing Eq.˜k-MFS corresponds to identifying a clique of size $k$, the decision version of the Maximum Clique problem. Intuitively, such cliques correspond to groups of components whose joint removal maximally increases the reconstruction error.
Computing the k-MFS
Since fidelity is monotonic, a natural heuristic selects the $k$ components with the highest singleton fidelities $s^{l}_{ci}$; we call this strategy: NAIVE. To compute the set of highest fidelity, the k-MFS, we identify conditions under which the NAIVE selection strategy is optimal. {restatable}theoremoptlabel Consider output channel $c$ in the $l^{th}$ layer of a network described in Section˜2.1. Let the $s^{l}_{ci}$ be defined according to Eq.˜3 and $S^{l\star}_{c}$ be defined according to Definition˜3.2. Let $\hat{S}^{l}_{c}=\{i\mid s^{l}_{ci}\geq s_{(k)}\}$ where $s_{(k)}$ is the $k^{th}$ largest value of $\bm{s}^{l}_{c}$. Assuming that there are no ties, $|\hat{S}^{l}_{c}|=k$. If $\mathbb{E}[\langle\bm{A}^{l}_{ci}(\mathrm{X}),\bm{A}^{l}_{cj}(\mathrm{X})\rangle]=0\ \ \forall i\neq j$, then $\hat{S}^{l}_{c}=S^{l\star}_{c}$.
Sketch.
Under the assumptions, the objective simplifies from quadratic to linear. See Section˜B.4. ∎
@@
Although the assumption of uncorrelated features rarely holds exactly in practice, it offers a sound theoretical justification for NAIVE HiFi selection. We demonstrate the practical effectiveness of NAIVE HiFi selection through our experiments in Section˜5.
@@ 4 Modifying Model Behavior using HiFi Sets
We now propose ModHiFi, a unified algorithmic framework for model modification using only distributional access. We apply this framework to two distinct tasks: structured pruning (ModHiFi-P) and classwise unlearning (ModHiFi-U). The central idea is to identify high-fidelity (HiFi) components and then modify them in a targeted manner using a unified algorithmic procedure, as shown in Algorithm˜1. The two tasks operate as duals: pruning retains the high-fidelity components necessary for general performance, while unlearning removes the high-fidelity components most discriminative for a specific target class. Additional details, including complexity and implementation specifics, are provided in Appendix˜D.
Structured Pruning
To address Eq.˜STRUCT-PRUNE, where the objective is to remove entire input channels (or features) that contribute minimally to the model’s predictive performance. In convolutional architectures, we identify and remove input channels across all layers that do not appear in the HiFi sets of any output channel of the residual-coupled layers. For CNNs, pruning is applied to the input channels of convolutional layers. For LLMs, we target the input features of the MLP down-projection matrices ($W^{D}$). After pruning, we compute the optimal compensation term $\delta^{\star}$ (derived in Section˜3.2) using the remaining weights. This step restores the fidelity of the layer output without requiring gradient-based fine-tuning.
@@
1:  Estimate Fidelity: Compute singleton scores $\mathbf{s}^{l}$ on $\mathcal{D}$ via Eq.˜3.
@@
The goal of Eq.˜UNLEARN is to erase the influence of a specific forget class. To perform unlearning, we first compute HiFi sets using only samples from the class we wish to forget. The components in these sets are then zeroed out, effectively erasing the influence of that class. This causes the model’s predictive performance on the forgotten class to degrade, without significantly impacting the performance of other classes.
@@
For vision models, the singleton fidelity score $\mathrm{FS}_{c}^{l}(\cdot)$ can be estimated efficiently using distributional access to the input data, i.e., synthetic samples. In practice, for vision models, we estimate the scalar coefficients $\alpha_{ci}^{l}$ directly via batched forward passes on synthetic samples. A large $\alpha_{ci}^{l}$ indicates a high-fidelity component. For LLMs, we develop a tractable Cholesky-based heuristic to estimate the score, providing details in Section˜D.2.
@@ 5 Experiments
    Effectiveness of using HiFi components for machine unlearning using ModHiFi-U. Is it possible to perform machine unlearning, as posed by Jia et al. [33], without finetuning? If so, how does ModHiFi-U compare to their method?
@@ 5.1 Details of the experimental setup
We conduct experiments on ResNet-50/101 [26], VGG19 [73], Swin-Transformer [46] and Llama-2-7B [80], benchmarking against relevant experiments from related literature [2, 48]. For vision tasks, we measure the classification accuracy, and for NLP tasks, we use EleutherAI’s lm-eval-harness [19].
@@
For CIFAR10/100 [37], we use synthetically generated images as detailed in Section˜C.3. We use Alpaca [78] (a synthetic dataset) and WikiText-2 [50] as calibration data for NLP tasks following related literature [2, 48]. We provide ablations to measure the impact of synthetic data quality in Section˜C.3.3.
@@
We discuss the compute platform, implementation details, and hyperparameters used for our experiments in Section˜C.6.
@@ 5.2 Existence of HiFi components: Exploring Q1
This empirical observation suggests that in trained models, only a small subset of components in each layer is responsible for the model’s prediction. This observation aligns with the success of structured pruning algorithms in constructing small subnetworks with high statistical performance.